NvidiaIntely (nvidiaintely):

Your uncle is making you a desk for you to take to college. One side of the desk is 48 inches and the diagonal length is 53.7 inches. What is the length of the other side of the desk? 12 inches 23.8 inches 17.4 inches 24 inches

NvidiaIntely (nvidiaintely):

How do I do this @mhchen?

NvidiaIntely (nvidiaintely):

@raffle_snaffle @Unavailabilityy @Will.H @wildemo

OpenStudy (mrnood):

@NvidiaIntely I am willing to help IF you are prepared to contribute (You haven't in other questions oyu posted) This involve the Pythagoras theorem and equation Can you sart by making a sketch (use drawing) of th etop of the desk, with th edimensions given in the question...

OpenStudy (mhchen):

So you use the pythagorean theorem. \[a^2 + b^2 = c^2\] where a and b are LEGS, and c is HYPOTENUSE So one side (LEG) is 48, and the diagonal (HYPOTENUSE) is 53.7 Plug it into the pythagorean theorem like this: \[48^2 + x^2 = 53.7^2\] and solve for x

OpenStudy (will.h):

There are so many people to help I suppose I'll just observe lol

NvidiaIntely (nvidiaintely):

How do I solve for x? And can you explain again what the pythagorean theorem is?

OpenStudy (will.h):

NV I thought you were a senior student

NvidiaIntely (nvidiaintely):

What's that?

OpenStudy (mrnood):

This is your usual approach - just keep asking until someone gives oyu the answer You MUST know the equation - it is what oyu are studying Or look it up

NvidiaIntely (nvidiaintely):

Can please tell me what it is?

OpenStudy (will.h):

You are high school student no?

NvidiaIntely (nvidiaintely):

Okay fine then! If no one wants to help me I will do it myself!

OpenStudy (mrnood):

Go and read your text or look it up on wiki it is a basic tool of trig and you should KNOW it , not need someone to give the answer

NvidiaIntely (nvidiaintely):

Thanks for trying @mhchen.

OpenStudy (will.h):

I think everyone been helping in the last few days. We would help you more if you get to learn how to solve it by yourself

OpenStudy (mrnood):

@mhchen ^ has DONE all the work except he simple solving for x

OpenStudy (mrnood):